# The Cooperative Network Architecture: Learning Structured Networks as Representation of Sensory Patterns

### v2

@@ 3 Computational Framework
Figure 2: The dynamics of a single area within the proposed CNA: The input image is fed into stage $S1$ to obtain feature activations. The feature activations are then processed by stage $S2$, together with activations from neurons within the same layer. In part A, two distinct matrices are used to model the forward connections from $S1$ to $S2$ and the recurrent connections within $S2$. In part B, the matrices and input signals are concatenated, resulting in one bigger weight matrix.
Our framework addresses the visual system, even though it could be applied to other modalities as well. In the primary visual cortex, images appear as two-dimensional arrays of activated neurons, and its neurons are sensitive to local texture elements within their receptive fields (Grill-Spector and Malach, 2004). To account for the homogeneity of image structure, it can be assumed that the connectivity has translational symmetry (‘convolutional’ structure) supporting equivariance, that is, the property that image shift gets reflected in activity shift. We use these insights to build a CNA with two stages, as shown in Figure 2(A). The CNA is based on an input layer representing observed images by simple feature detectors and two layers, Stage 1 (denoted as $S1$) and Stage 2 (denoted as $S2$).
Feature Extraction.
The system’s input is a retinal 2D image represented by variables $\boldsymbol{x}_{k,i}$ (gray box in Figure 2), where $k\in(1,...,C)$ denotes the channel index of the image ($C=1$ for grayscale, $C=3$ for RGB), and the composite index $i=(h,w)\in(H\times W)$ stands for vertical and horizontal position of pixels within the input image.
Such an image is used as afferent input to the feature extraction neurons in $S1$ and first increases their ‘membrane potential’ $\boldsymbol{a}^{(S1)}\in\mathbb{R}^{C^{(S1)}\times H\times W}$ and potentially triggers their binary activation state $\boldsymbol{y}^{(S1)}\in\{0,1\}^{C^{(S1)}\times H\times W}$ (green box in Figure 2). The membrane potential $\boldsymbol{a}^{(S1)}$ is calculated by applying a convolutional operation (LeCun et al., 1989) (yellow box $S1$ in Figure 2):
$\boldsymbol{a}^{(S1)}_{c,j}=\sum_{k\in C}\sum_{i\in l_{j}}\boldsymbol{W}^{(S1)}_{c,k,j-i}\cdot\boldsymbol{x}_{k,i}$ (1)
Here, $c\in C^{(S1)}$ denotes the feature channel, the compound index $j$, expanded as $j=\{h,w\}$, designates again spatial position $h\in H$, $w\in W$, and the compound index $i=\{h,w\}$ is running through all $(h,w)$-pairs in the two-dimensional range $l_{j}=(h,....,h+h^{(W1)};w,....,w+w^{(W1)})$ where $h^{(W1)}$, $w^{(W1)}$ denote the size of the convolutional filter. The size of the convolutional kernel used for feature extraction is $\boldsymbol{W}^{(S1)}\in\mathbb{R}^{C^{(S1)}\times C\times h^{(W1)}\times w^{(W1)}}$.
Feature Binarization.
The membrane potential $\boldsymbol{a}^{(S1)}$, together with a bias parameter $b^{(S1)}$, determines the binary output of the neurons in $S1$, denoted as variable $\boldsymbol{y}_{c,j}^{(S1)}\in\{0,1\}^{C^{(S1)}\times H\times W}$, where $1$ stands for firing, 00 for silence. The binary neuron model with the input-output function $\mathrm{B}:\boldsymbol{a},b\rightarrow\boldsymbol{y}$ is defined as
$\boldsymbol{y}_{c,j}^{(S1)}=\mathrm{B}(\boldsymbol{a}^{(S1)}_{c,j}-b^{(S1)})=\begin{cases}1,&\text{ if }\boldsymbol{a}^{(S1)}_{c,j}-b^{(S1)}>0\\ 0,&\text{otherwise}\end{cases}$ (2)
with the same bias parameter $b^{(S1)}$ for all neurons acting as a firing threshold (neurons fire if the membrane potential is above this threshold).
Dynamics.
Stage $S2$ builds net fragments by allowing active neurons to support recurrently connected neurons over T time steps (depicted as recurrent connection in Figure 2). A sample image is introduced into the system at time step $t=0$ to activate neurons and remains unchanged until time step $t=T$. Throughout this period, neuronal activity undergoes sparsification due to growing inhibition (see below). Through the sparsification process, the system forms an object representation by leaving active a set of neurons each of which is supported by active neighbors, that is, a net.
Short-Range recurrent Connections.
Although typical convolutional neural networks (Fukushima, 1980; Waibel et al., 1987; LeCun et al., 1989) apply convolution operations to establish feedforward connections between different layers, we employ them here to implement short-range recurrent connections within the same layer, their symmetry permitting a pattern to manifest itself at any spatial location based on corresponding net fragments. Given recurrent connections, the neural activity of the previous time step $t-1$ is accessed to calculate a neuron’s activity at time $t$. In contrast to recurrent networks like LSTMs (Hochreiter and Schmidhuber, 1997), where a new token is input at each time step, we iterate over the same input while updating the internal state to project out a net representing a given sensory signal. Importantly, the internal state evolves continuously, even when processing the same input pattern, driven by dynamic changes in the support between neurons along recurrent connections.
The membrane potentials $\boldsymbol{a}^{(S2)}_{c,j}$ of the second stage neurons get excited by the (constant) signals of the previous stage $S1$ and iteratively by recurrent excitation from within $S2$. Before digitization, they are modified by saturation, normalization, and inhibition (yellow box $S2$ in Figure 2). Before those modifications, its raw form is computed as:
@@
$\boldsymbol{a}^{\prime(S2)}_{c,j}[t]=\underbrace{\left(\sum_{k\in C^{(S1)}}\sum_{i\in l_{j}}\boldsymbol{W}^{(F)}_{c,k,j-i}\cdot\boldsymbol{y}_{k,i}^{(S1)}\right)}_{\textit{Forward Connection}}+\overbrace{\left(\sum_{k\in C^{(S2)}}\sum_{i\in l_{j}}\boldsymbol{W}^{(L)}_{c,k,j-i}\cdot\boldsymbol{y}_{k,i}^{(S2)}[t-1]\right)}^{\textit{Recurrent Connection}}.$ (3)
Similar to $S1$, $c\in C^{(S2)}$ denotes the feature channel of $S2$, the compound index $j$ expands to $j=\{w,h\}$, $w\in W$, $h\in H$, and the compound index $i$ is running in the range $l_{j}=(h,....,h+h^{(W2)};w,....,w+w^{(W2)})$, with $h^{(W2)}$ and $w^{(W2)}$ being the size of the convolution kernel.
The forward connections are denoted as $\boldsymbol{W}^{(F)}\in\mathbb{R}^{C^{(S2)}\times C^{(S1)}\times h^{(W2)}\times w^{(W2)}}$ and the recurrent connections as $\boldsymbol{W}^{(L)}\in\mathbb{R}^{C^{(S2)}\times C^{(S2)}\times h^{(W2)}\times w^{(W2)}}$, where $C^{(S2)}$ is the number of channels in $S2$, including the set of all feature channels and its alternative channels, i.e. $C^{(S2)}=n_{a}\cdot C^{(S1)}$ where $n_{a}$ denotes the number of alternative channels per feature type (the alternative channels handle the alternative neurons that are required to prevent cross-talk between fragments). In the case of the forward connections, the kernel size determines which of the neurons $\boldsymbol{y}^{(S1)}$ in stage 1 can be connected to a neuron $\boldsymbol{y}^{(S2)}$ in stage 2 (red box in Figure 2), while the kernel size of the recurrent connections determines which neurons can be connected within stage 2. The kernel $\boldsymbol{W}^{(F)}$ is initialized such that the sets of equivalent neurons in $S2$ copy the activity (and with that the feature type) of the corresponding neuron in $S1$, and $\boldsymbol{W}^{(L)}$ is initialized with zeros except for self-coupling of neurons, with value 1 (details in Appendix A). The weights are updated using Hebbian Learning (details in Appendix B).
For the first time step $t=0$, when $\boldsymbol{y}_{i}^{(S2)}$ is undefined, we initialize the $S2$ neurons with zeros, i.e., the neurons are silent and do not provide recurrent support:
$\boldsymbol{y}_{i}^{(S2)}[t=0]=0$ (4)
Single Weight-Kernel Implementation.
This entire process is visually depicted in the upper part (A) of Figure 2. For simplicity in our algorithm, and since we use the same kernel size for the forward and recurrent connections, we stack the two kernels $\boldsymbol{W}^{(F)}$ and $\boldsymbol{W}^{(L)}$ as $\boldsymbol{W}^{(S2)}=(\boldsymbol{W}^{(F)};\boldsymbol{W}^{(L)})$ as well as the inputs $\boldsymbol{y}_{k,i}^{(S1)}$ and $\boldsymbol{y}_{k,i}^{(S2)}[t-1]$ as shown in Figure 2(B), leading to the equivalent formulation:
$\boldsymbol{a}^{\prime(S2)}_{c,j}[t]=\sum_{k\in C^{(S1)}}\sum_{i\in l_{j}}\boldsymbol{W}^{(S2)}_{c,k,j-i}\cdot\left(\boldsymbol{y}_{k,i}^{(S1)};\boldsymbol{y}_{k,i}^{(S2)}[t-1]\right)$ (5)
The net fragments are implicitly contained in this processing as the active neurons whose activity state is stored in $\boldsymbol{y}^{(S1)}$ and their connection to other neurons.
Inhibition.
The raw neural excitation $\boldsymbol{a}^{\prime}$ computed in eq. 3 (or eq. 5) is to be subjected to saturation, inhibition, and normalization so it can be binarized using a pre-defined threshold. These steps of processing the raw membrane potential $\boldsymbol{a}^{\prime}$ of neurons proved necessary in our simulation experiments to achieve the qualitative function we aimed for. The details to obtain the normalized activations $\boldsymbol{a}^{\text{norm}}$ in the range $(0,1)$ are described in Appendix C. The differences between activation levels of neurons are accentuated by setting
$\boldsymbol{a}^{(S2)}_{c,j}[t]=\max\left(\left(\boldsymbol{a}^{\text{norm }(S2)}_{c,j}[t]\right)^{\gamma},0\right),$ (6)
The coefficient $\gamma$ regulates the inhibition strength and increases with the time-steps of the inner loop according to $\gamma=1.2+0.2t$, making it harder for neurons to fire (as $\boldsymbol{a}^{\text{norm }(S2)}\leq 1$). The membrane potential is higher for neurons that are part of coherent fragments (valid patterns) as they receive recurrent support, while invalid activations are less supported. While the inhibition coefficient increases over time, many neurons eventually get suppressed, leaving only those active that support each other and form a consistent net. Finally, the membrane potential is converted to a binary activation, neurons becoming active when their inhibited membrane potential is above the activation bias $b^{(S2)}$, which serves to control the firing rate of neurons:
$\boldsymbol{y}_{c,j}^{(S2)}[t]=\mathrm{B}\left(\boldsymbol{a}^{(S2)}_{c,j}[t]-b^{(S2)}\right)$ (7)
The final output $\boldsymbol{y}^{(S2)}[t=T]$ is the recurrent connection-supported neuronal activation in which, in contrast to the initial activation, unsupported activations are suppressed. Alternative neurons are connected by distinct channels within the convolutional matrix and engage in competition, with only one alternative channel gaining activation at each spatial location while the remaining neurons are suppressed (set to zero). This selection process process is described in the Appendix D.
Relation to Nets and Fragments.
Net fragments are learned through the strengthening of recurrent connections $\boldsymbol{W}^{(L)}$, facilitated by Hebbian plasticity. Neurons within a fragment are interconnected, providing mutual support that enhances their membrane potentials. By shifting $\boldsymbol{W}^{(L)}$ as convolutional kernels across the input, fragments can be detected at all spatial locations, resulting in numerous overlapping fragments. Coherent fragments are interconnected to form a larger structure, a net.
While multiple nets may exist, projecting out the most coherent one is crucial for robust object representation. This is achieved by progressively increasing inhibition over time. As inhibition intensifies, neurons with insufficient support (e.g., only a few instead of many neurons within one net fragment were triggered) are deactivated [^2] Neuron activations can be interpreted as hypotheses that require confirmation/support from other neurons with compatible hypotheses - otherwise, they are deactivated.. This deactivation can propagate, leading to the elimination of entire fragments and leaving a single coherent net to represent the input structure.
In our experiments (see below), we observe that, initially, neurons operate independently due to the absence of established recurrent connections, with a single neuron having a membrane potential of $1$ being sufficient for activation. However, after training, the average membrane potential rises to approximately $12$, reflecting robust support from neurons within a fragment. As a result, inconsistent neurons can no longer activate (and neurons with a membrane potential of $1$ are switched off), ensuring a more stable and coherent representation.

### v3

@@ 3 Cooperative Network Architecture
Figure 2: The dynamics within the proposed CNA: The input image is fed into stage $S1$ to obtain feature activations. The feature activations are then processed by stage $S2$, together with activations from neurons within the same layer.
In this section, we present a concrete instantiation of the proposed framework, specifically applied to the visual domain – although the framework is general and may be extended to other sensory modalities in the future. We refer to this implementation as the Cooperative Network Architecture (CNA). To realize our framework, we adopt a standard convolutional neural network (CNN) architecture (LeCun et al., 1989).
The CNA comprises two processing stages, depicted as yellow squares in Figure 2. The stage 1 ($S1$) performs fixed (i.e., not learned) feature extraction directly from the image input. The stage 2 ($S2$) implements the core principles of the described computational framework by learning net fragments.
The network dynamics unfold over discrete time steps $t=[0,\dots,T]$. Since the feature extraction stage ($S1$) employs fixed filters, it yields a time-invariant activation pattern for a given input. In contrast, the second stage ($S2$) exhibits dynamic temporal evolution, where internal representations are iteratively refined at each time step ($t\to t+1$). Specifically, we use a progressively increasing global attenuation signal that only allows neurons with a high membrane potential to remain active. During this dynamic processing, a neuron might be deactivated, whereby another neuron’s membrane potential is decreased, triggering a chain reaction that can lead to the deactivation of incoherent nets.
While these refinements occur within the processing cycle of a single sample ($[0,\dots,T]$), Hebbian learning adapts the weights (connections $w_{i,j}$ within net fragments) on longer timescales, wherein synaptic weight modifications occur between different sample-processing episodes, leading to structural plasticity in the network. The following sections discuss these two processing stages.
3.1 Stage 1: Feature Extraction
Stage 1 is designed to model the sensory input by extracting initial neural activations from the raw image data. This extraction is essential due to the continuous nature of typical image inputs (e.g., grayscale or RGB), whereas the computational framework employed in this work operates on binary activations. Consequently, this stage serves as a preprocessing step, enabling the transformation of continuous-valued images into a suitable binary representation.
The input image is denoted as $\boldsymbol{x}\in\mathbb{R}^{C,H,W}$ (illustrated by the gray box in Figure 2), where $C$ indicates the number of image channels ($1$ for grayscale, $3$ for RGB), while $H,W$ stand for the height and width of the image.
Activation Computation.
To extract initial activation patterns, we use fixed, non-trainable filters (details see Appendix A). These filters are implemented as standard convolutional kernels, used to obtain the binary activation pattern $\boldsymbol{y}^{(S1)}\in\{0,1\}^{C^{(S1)},H,W}$ as follows:
$\boldsymbol{y}^{(S1)}_{c_{\text{out}}}=\Theta\left(\sum_{c_{\text{in}}=0}^{C}\boldsymbol{W}^{(S1)}_{c_{\text{out}},c_{\text{in}}}\star\boldsymbol{x}_{c_{\text{in}}}\right)$ (12)
Here, $\star$ denotes the convolution operator between the filters $\boldsymbol{W}^{(S1)}\in\mathbb{R}^{C^{(S1)},C,h,w}$ and the input image $\boldsymbol{x}$, while the parameters $h,w$ define the width and height of the convolutional filters, i.e., the spatial neighborhood from where the features are extracted. To binarize the activations, we use the heaviside function $\Theta$, which sets a neuron to one if its input is larger than zero, and to zero otherwise.
$y_{i}=\Theta(a_{i})=\begin{cases}1\quad\text{if }a_{i}>0\\ 0\quad\text{otherwise}\end{cases}$ (13)
Note that the width and height between input and output activations remain identical, but the number of channels changes from $C$ to $C^{(S1)}$, allowing the extraction of multiple features at each spatial location.
3.2 Stage 2: Net Fragments
$S2$ constructs net fragments from the binary feature activations $\boldsymbol{y}^{(S1)}$ of $S1$ and integrates them into coherent, spatially extended nets over time. To ensure stability, single neurons or small groups of neurons are suppressed through an attenuation mechanism, ensuring that only large-scale, consistent activation patterns persist.
Network Topology.
Stage 2 uses convolutional layers to implement the connections $w_{j,i}$ within net fragments $n_{j}\in\mathcal{F}(n_{i})$. Conceptually, this stage comprises a three-dimensional array of neurons with dimensions $C^{(S2)}\times H\times W$, where $C^{(S2)}$ denotes the number of feature channels in $S2$, and $H$ and $W$ correspond to the spatial dimensions of the input image.
The spatial distance between two neurons $n_{i}$ and $n_{j}$ is defined using the maximum norm (also known as Chebyshev distance), where adjacent neurons have a unit distance:
$d(n_{i},n_{j})=\max\left(|h_{i}-h_{j}|,|w_{i}-w_{j}|\right)$ (14)
Each neuron $n^{(S2)}_{i}$ in stage 2 is connected in two ways: (1) Feedforward connections to the corresponding neuron $n^{(S1)}_{i}$ in stage 1 and to its local neighborhood $Nr(n^{(S1)}_{i})$, and (2) lateral connections to neighboring neurons within stage 2 itself, denoted as:
$Nr\left(n^{(S2)}_{i}\right)=Nr\left(n^{(S1)}_{i}\right)\cup\left\{n_{j}^{(S2)}\mid d(n_{i}^{(S2)},n_{j}^{(S2)})\leq r_{2}\right\}$ (15)
where $r_{2}$ defines the radius of lateral connections. The convolutional architecture realizes these neighborhood structures via convolutional filters, where the filter size directly determines the radius of both feedforward and lateral connections. Using convolutional filters not only implements the connectivity restrictions to local neighborhoods but also allows detection of the same pattern at various positions, enabling learning of net fragments independent of position.
The forward connections from $S1$ to $S2$, denoted as $\boldsymbol{W}^{(F)}$, are necessary to activate cells within $S2$. It can be interpreted as a small group of neurons in $S1$ (based on the filter size of $\boldsymbol{W}^{(F)}$) supporting a single neuron within $S2$. In contrast, the lateral recurrent connections within $S2$, denoted as $\boldsymbol{W}^{(L)}$, are responsible for building the overlapping net fragments within $S2$, providing support amongst the neurons within $S2$. It is important to note that the lateral connection matrix $\boldsymbol{W}^{(L)}$ does not implement recurrent processing in the traditional sense of sequential token-wise updates, as found in RNN-based language models. Instead, $\boldsymbol{W}^{(L)}$ iterates over the same neuronal state vector $\boldsymbol{y}^{(S2)}$. Specifically, the lateral connections $\boldsymbol{W}^{(L)}$ calculate the state updates of $\boldsymbol{y}^{(S2)}_{t\rightarrow t+1}$ within an iterative refinement process in which individual neuronal activations ${y}_{i}^{(S2)}\in\boldsymbol{y}^{(S2)}$ may flip due to changes in support and attenuation.
Membrane Potential.
To determine the support a single neuron $n_{i,t}$ receives from the net fragment $\mathcal{F}_{t-1}(n_{i,t-1})$, we first calculate the membrane potential $\mathcal{M}$ by applying convolutional operations on the forward and lateral connections and then taking the sum of these operations:
@@ 3.2 Stage 2: Net Fragments
$\boldsymbol{a}^{\mathcal{M}(S2)}_{c_{\text{out}},t}=\underbrace{\left(\sum_{c_{\text{in}}=0}^{C^{(S1)}}\boldsymbol{W}^{(F)}_{c_{\text{out}},c_{\text{in}}}\star\boldsymbol{y}_{c_{\text{in}}}^{(S1)}\right)}_{\textit{Forward Connection}}+\overbrace{\left(\sum_{c_{\text{in}}=0}^{C^{(S2)}}\boldsymbol{W}^{(L)}_{c_{\text{out}},c_{\text{in}}}\star\boldsymbol{y}_{c_{\text{in}},t-1}^{(S2)}\right)}^{\textit{Recurrent Connection}}.$ (16)
Here, the forward connections have the shape $\boldsymbol{W}^{(F)}\in\mathbb{R}^{C^{(S2)},C^{(S1)},h,w}$ and the recurrent connections have the shape $\boldsymbol{W}^{(L)}\in\mathbb{R}^{C^{(S2)},C^{(S2)},h,w}$, where $C^{(S2)}$ is the number of channels in $S2$. Therefore, the membrane potential $\mathcal{M}$ at time $t$ is increased by the learned amount $w_{j,i}$ from all connected and active neurons ($y_{j,t-1}=1$). Note that $\mathcal{F}_{0}(n_{i,0})$ only consists of neurons of $S1$ since all neurons in $S2$ are initialized to silent state ($y_{i,0}=0$).
Higher support, indicated by larger $a^{\mathcal{M}(S2)}_{i}$, suggests a more likely pattern, as neurons encoding frequent patterns gain reinforcement not only from feedforward connections but also from lateral connections. Before we discuss how the weight matrices $\boldsymbol{W}^{(F)},\boldsymbol{W}^{(L)}$ are initialized and trained, we describe how the membrane potential $\boldsymbol{a}^{\mathcal{M}(S2)}$ undergoes attenuation and is transformed into $\boldsymbol{y}^{(S2)}$.
Attenuation.
The membrane potential is processed by an activation function $\mathcal{A}\left(\boldsymbol{a}^{\mathcal{M}},\mathcal{E}\right)$. In our implementation, this function first normalizes the potential $\boldsymbol{a}^{\mathcal{M}}\rightarrow\boldsymbol{a}^{\text{norm}}$ to conform it into the range $[0,1]$ (details see Appendix B), and then applies a globally increasing attenuation pressure $\boldsymbol{a}^{\text{norm}}\rightarrow\boldsymbol{a}$ to ensure that only neurons receiving substantial support remain active. Specifically, attenuation is implemented as:
$\boldsymbol{a}^{(S2)}_{c,t}=\left(\boldsymbol{a}^{\text{norm }(S2)}_{c,t}\right)^{\gamma},$ (17)
where $\gamma$ regulates the attenuation strength and increases over time according to $\gamma=\alpha+\beta t$ (when not stated differently, we set $\alpha=1.2$ and $\beta=0.2$), making activation progressively harder. This process leads to a sparsification of activations that occurs through both direct attenuation and reduced support from previously silenced neurons .
The final output is then binarized using the heaviside function $\Theta$:
$\boldsymbol{y}_{c,t}^{(S2)}=\Theta\left(\boldsymbol{a}^{(S2)}_{c,t}-b^{(S2)}\right)$ (18)
The activation at the last timestep $\boldsymbol{y}^{(S2)}_{t=T}$ represents a recurrently refined activation pattern where unsupported neurons have been suppressed, yielding a coherent representation.
Alternative Neurons.
A single neuron that encodes a specific feature can be activated in multiple contexts. For example, a short straight-line segment may appear in various contexts – such as in different line drawings or digits. In our framework, each context is represented by a net, whereas individual features correspond to the activity of specific neurons.
To achieve context disentanglement, it is important to allocate sufficient representational capacity such that the same feature (e.g., a line segment) can be utilized across distinct contexts (e.g., different line drawings) without interference. This requires that a feature-representing neuron is capable of participating in multiple, context-specific net fragments. Without this capacity, a neuron would be forced to average over competing contextual associations, resulting in a non-specific or “blurry” representation.
Formally, a neuron representing a feature may be part of multiple net fragments: $n_{i}\in\mathcal{F}(n_{j})$ and $n_{i}\in\mathcal{F}(n_{k})$ where $n_{j}\neq n_{k}$). This implies that the neuron must form distinct lateral connections $w_{i,j}$ and $w_{i,k}$, which should not be concurrently active.
To address this requirement, we introduce “alternative neurons”. Each alternative version of a neuron is activated by the same feature from $S1$ but has a distinct set of lateral connections in $S2$ to encode different contexts. This is implemented by increasing the channel dimension $C^{(S2)}$ of $\boldsymbol{W}^{(L)}$ by a factor $\kappa$, i.e.,
$C^{(S2)}:=C^{(S2)}\cdot\kappa$ (19)
This expansion provides capacity for $\kappa$ alternative version of a neuron. Consequently, these alternative neurons receive identical bottom-up input from $S1$ while differing in the lateral support they receive in $S2$. A winner-take-all (WTA) competition mechanism ensures that each alternative neuron specializes in a distinct context, with only one being activated in any given instance.
[…]
At initialization, neuron selection is random. However, after Hebbian learning updates, the alternative neurons progressively specialize, leading to distinct and structured fragment representations. Over time, the most correlated alternative neuron is chosen for each input location, ensuring robust feature specialization across different contexts.
Weight Initialization.
The forward connections $\boldsymbol{W}^{(F)}$ are initialized so that the activations of the $S1$ neurons are initially copied into all alternative $S2$ neurons in the same position. This is done by setting the connections at the center of the kernel (at position $(h/2,w/2)$) to $1$ if $c_{in}=\lceil c_{out}/\kappa\rceil$, where $c_{in}$ denotes the index of the input channel, $c_{out}$ the index of the output channel, and $\kappa$ the number of alternative neurons. The recurrent connections $\boldsymbol{W}^{(L)}$ are initialized with zeros except for the self-coupling of neurons (recurrent connection to itself), which is set to $1$.
Learning of Weights.
Support within net fragments is learned via Hebbian plasticity (Hebb, 1949), so that frequently occurring patterns that lead to correlated activation patterns between neurons $n_{i}\sim n_{j}$, are “stored” in the system and receive more support:
$\boldsymbol{W}^{(S2)}:=\min\left(\max\left(\boldsymbol{W}^{(S2)}+\eta\cdot\boldsymbol{\rho}_{avg},0\right),1\right)$ (20)
Here, $\eta$ represents the learning rate and $\boldsymbol{W}^{(S2)}$ stands for both $\boldsymbol{W}^{(F)}$ and $\boldsymbol{W}^{(L)}$. Note that the same convolutional kernel is applied at various spatial positions. The decision to update this connection is based on the average correlation $\boldsymbol{\rho}_{avg}$ between all input and output neurons it connects (across all spatial positions). If the resulting average is positive, the corresponding connection increases as it connects more simultaneously active neurons than neurons that fire disjointly. Conversely, if the average correlation is negative, indicating more disjoint firing, the connection strength is reduced. Learning is applied after processing a sample (after $T$ timesteps) to improve stability.